Chapter 1

Contribution Title

Name of First Author and Name of Second Author

Abstract Each chapter should be preceded by an abstract (10–15 lines long) that summarizes the content. The abstract will appear online at www.SpringerLink.com and be available with unrestricted access. This allows unregistered users to read the abstract as a teaser for the complete chapter. As a general rule the abstracts will not appear in the printed version of your book unless it is the style of your particular book or that of the series to which your book belongs.

Please use the ’starred’ version of the new Springer abstract command for typesetting the text of the online abstracts (cf. source file of this chapter template abstract) and include them with the source files of your manuscript. Use the plain abstract command if the abstract is also to appear in the printed version of the book.

1.1 Section Heading

Use the template chapter.tex together with the Springer document class SVMono (monograph-type books) or SVMult (edited books) to style the various elements of your chapter content in the Springer layout.

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations. And please note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

Name of First Author Name, Address of Institute, e-mail: name@email.address Name of Second Author Name, Address of Institute e-mail: name@email.address

1.2 Section Heading

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations.

Please note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

Use the standard equation environment to typeset your equations, e.g.

a × b = c ,  (1.1)

however, for multiline equations we recommend to use the eqnarray environment^{1}.

a × b = c
a· b = c  (1.2)

1.2.1 Subsection Heading

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations as has already been described in Sect. 1.2.

Please do not use quotation marks when quoting texts! Simply use the quotation environment – it will automatically render Springer’s preferred layout.

1.2.1.1 Subsubsection Heading

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations as has already been described in Sect. 1.2.1, see also Fig. 1.1^{2}

Please note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

1 In physics texts please activate the class option vecphys to depict your vectors in boldface-italic type - as is customary for a wide range of physical subjects 2 If you copy text passages, figures, or tables from other works, you must obtain permission from the copyright holder (usually the original publisher). Please enclose the signed permission with the manuscript. The sources must be acknowledged either in the captions, as footnotes or in a separate section of the book.

Paragraph Heading

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations as has already been described in Sect. 1.2.

Please note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

For typesetting numbered lists we recommend to use the enumerate environment – it will automatically render Springer’s preferred layout.

1. Livelihood and survival mobility are oftentimes coutcomes of uneven socioeconomic development.

a. Livelihood and survival mobility are oftentimes coutcomes of uneven socioe-

conomic development. b. Livelihood and survival mobility are oftentimes coutcomes of uneven socioe-

conomic development.

2. Livelihood and survival mobility are oftentimes coutcomes of uneven socioeconomic development.

Subparagraph Heading

In order to avoid simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Use the L^{A}TEX automatism for all your cross-references and citations as has already been described in Sect. 1.2, see also Fig. 1.2.

For unnumbered list we recommend to use the itemize environment – it will automatically render Springer’s preferred layout.

• Livelihood and survival mobility are oftentimes coutcomes of uneven socioeconomic development, cf. Table 1.1.

Fig. 1.1 If the width of the figure is less than 7.8 cm use

the sidecapion command

to flush the caption on the left side of the page. If the figure is positioned at the top of the page, align the sidecaption with the top of the figure – to achieve this you simply need to use the optional argument

[t] with the sidecaption

command Fig. 1.2 If the width of the figure is less than 7.8 cm use

the sidecapion command

to flush the caption on the left side of the page. If the figure is positioned at the top of the page, align the sidecaption with the top of the figure – to achieve this you simply need to use the optional argument

[t] with the sidecaption

command

– Livelihood and survival mobility are oftentimes coutcomes of uneven socioeconomic development.

– Livelihood and survival mobility are oftentimes coutcomes of uneven socioeconomic development.

• Livelihood and survival mobility are oftentimes coutcomes of uneven socioeconomic development.

Run-in Heading Boldface Version Use the L^{A}TEX automatism for all your crossreferences and citations as has already been described in Sect. 1.2.

Run-in Heading Italic Version Use the L^{A}TEX automatism for all your cross-refer-ences and citations as has already been described in Sect. 1.2.

Table 1.1 Please write your table caption here

Classes  Subclass  Length  Action Mechanism
Translation  mRNA^{a}  22 (19–25)  Translation repression, mRNA cleavage
Translation  mRNA cleavage 21  mRNA cleavage
Translation  mRNA  21–22  mRNA cleavage
Translation  mRNA  24–26  Histone and DNA Modification

a Table foot note (with superscript)

1.3 Section Heading

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations as has already been described in Sect. 1.2.

Please note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

If you want to list definitions or the like we recommend to use the Springerenhanced description environment – it will automatically render Springer’s preferred layout.

Type 1 That addresses central themes pertainng to migration, health, and disease. In Sect. 1.1, Wilson discusses the role of human migration in infectious disease distributions and patterns.

Type 2 That addresses central themes pertainng to migration, health, and disease. In Sect. 1.2.1, Wilson discusses the role of human migration in infectious disease distributions and patterns.

1.3.1 Subsection Heading

In order to avoid simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Use the L^{A}TEX automatism for all your cross-references and citations citations as has already been described in Sect. 1.2.

Please note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

If you want to emphasize complete paragraphs of texts we recommend to use the newly defined Springer class option graybox and the newly defined environment svgraybox. This will produce a 15 percent screened box ’behind’ your text.

If you want to emphasize complete paragraphs of texts we recommend to use the newly defined Springer class option and environment svgraybox. This will produce a 15 percent screened box ’behind’ your text.

1.3.1.1 Subsubsection Heading

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations as has already been described in Sect. 1.2.

Please note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

Theorem 1. Theorem text goes here.

Definition 1. Definition text goes here.

Proof. Proof text goes here. ⊓ ⊔

Paragraph Heading

Instead of simply listing headings of different levels we recommend to let every heading be followed by at least a short passage of text. Further on please use the L^{A}TEX automatism for all your cross-references and citations as has already been described in Sect. 1.2.

Note that the first line of text that follows a heading is not indented, whereas the first lines of all subsequent paragraphs are.

Theorem 2. Theorem text goes here.

Definition 2. Definition text goes here.

Proof. Proof text goes here. ⊓ ⊔

Acknowledgements If you want to include acknowledgments of assistance and the like at the end of an individual chapter please use the acknowledgement environment – it will automatically render Springer’s preferred layout.

Appendix

When placed at the end of a chapter or contribution (as opposed to at the end of the book), the numbering of tables, figures, and equations in the appendix section continues on from that in the main text. Hence please do not use the appendix command when writing an appendix at the end of your chapter or contribution. If there is only one the appendix is designated “Appendix”, or “Appendix 1”, or “Appendix 2”, etc. if there is more than one.

a × b = c  (1.3)

References

References may be cited in the text either by number (preferred) or by author/year.^{3} The reference list should ideally be sorted in alphabetical order – even if reference numbers are used for the their citation in the text. If there are several works by the same author, the following order should be used:

3 Make sure that all references from the list are cited in the text. Those not cited should be moved to a separate Further Reading section or chapter.

1. all works by the author alone, ordered chronologically by year of publication

2. all works by the author with a coauthor, ordered alphabetically by coauthor

3. all works by the author with several coauthors, ordered chronologically by year of publication.

The styling of references^{4} depends on the subject of your book:

• The two recommended styles for references in books on mathematical, physical, statistical and computer sciences are depicted in [1, 2, 3, 4, 5] and [6, 7, 8, 9, 10].

• Examples of the most commonly used reference style in books on Psychology, Social Sciences are [11, 12, 13, 14, 15].

• Examples for references in books on Humanities, Linguistics, Philosophy are [16, 17, 18, 19, 20].

• Examples of the basic Springer style used in publications on a wide range of subjects such as Computer Science, Economics, Engineering, Geosciences, Life Sciences, Medicine, Biomedicine are [21, 22, 24, 23, 25].

1. Broy, M.: Software engineering — from auxiliary to key technologies. In: Broy, M., Dener, E. (eds.) Software Pioneers, pp. 10-13. Springer, Heidelberg (2002)

2. Dod, J.: Effective substances. In: The Dictionary of Substances and Their Effects. Royal Society of Chemistry (1999) Available via DIALOG. http://www.rsc.org/dose/title of subordinate document. Cited 15 Jan 1999

3. Geddes, K.O., Czapor, S.R., Labahn, G.: Algorithms for Computer Algebra. Kluwer, Boston (1992)

4. Hamburger, C.: Quasimonotonicity, regularity and duality for nonlinear systems of partial differential equations. Ann. Mat. Pura. Appl. 169, 321–354 (1995)

5. Slifka, M.K., Whitton, J.L.: Clinical implications of dysregulated cytokine production. J. Mol. Med. (2000) doi: 10.1007/s001090000086

6. J. Dod, in The Dictionary of Substances and Their Effects, Royal Society of Chemistry. (Available via DIALOG, 1999), http://www.rsc.org/dose/title of subordinate document. Cited 15 Jan 1999

7. H. Ibach, H. Lüth, Solid-State Physics, 2nd edn. (Springer, New York, 1996), pp. 45-56

8. S. Preuss, A. Demchuk Jr., M. Stuke, Appl. Phys. A 61

9. M.K. Slifka, J.L. Whitton, J. Mol. Med., doi: 10.1007/s001090000086

10. S.E. Smith, in Neuromuscular Junction, ed. by E. Zaimis. Handbook of Experimental Pharmacology, vol 42 (Springer, Heidelberg, 1976), p. 593

11. Calfee, R. C., & Valencia, R. R. (1991). APA guide to preparing manuscripts for journal publication. Washington, DC: American Psychological Association.

12. Dod, J. (1999). Effective substances. In: The dictionary of substances and their effects. Royal Society of Chemistry. Available via DIALOG. http://www.rsc.org/dose/Effective substances. Cited 15 Jan 1999.

13. Harris, M., Karper, E., Stacks, G., Hoffman, D., DeNiro, R., Cruz, P., et al. (2001). Writing labs and the Hollywood connection. J Film Writing, 44(3), 213–245.

14. O’Neil, J. M., & Egan, J. (1992). Men’s and women’s gender role journeys: Metaphor for healing, transition, and transformation. In B. R. Wainrig (Ed.), Gender issues across the life cycle (pp. 107–123). New York: Springer.

4 Always use the standard abbreviation of a journal’s name according to the ISSN List of Title Word Abbreviations, see http://www.issn.org/en/node/344 15. Kreger, M., Brindis, C.D., Manuel, D.M., Sassoubre, L. (2007). Lessons learned in systems

change initiatives: benchmarks and indicators. American Journal of Community Psychology, doi: 10.1007/s10464-007-9108-14.

16. Alber John, Daniel C. O’Connell, and Sabine Kowal. 2002. Personal perspective in TV interviews. Pragmatics 12:257–271

17. Cameron, Deborah. 1997. Theoretical debates in feminist linguistics: Questions of sex and gender. In Gender and discourse, ed. Ruth Wodak, 99–119. London: Sage Publications.

18. Cameron, Deborah. 1985. Feminism and linguistic theory. New York: St. Martin’s Press.

19. Dod, Jake. 1999. Effective substances. In: The dictionary of substances and their effects. Royal Society of Chemistry. Available via DIALOG. http://www.rsc.org/dose/title of subordinate document. Cited 15 Jan 1999

20. Suleiman, Camelia, Daniel C. OConnell, and Sabine Kowal. 2002. ‘If you and I, if we, in this later day, lose that sacred fire...’: Perspective in political interviews. Journal of Psycholinguistic Research. doi: 10.1023/A:1015592129296.

21. Brown B, Aaron M (2001) The politics of nature. In: Smith J (ed) The rise of modern genomics, 3rd edn. Wiley, New York

22. Dod J (1999) Effective Substances. In: The dictionary of substances and their effects. Royal Society of Chemistry. Available via DIALOG. http://www.rsc.org/dose/title of subordinate document. Cited 15 Jan 1999

23. Slifka MK, Whitton JL (2000) Clinical implications of dysregulated cytokine production. J Mol Med, doi: 10.1007/s001090000086

24. Smith J, Jones M Jr, Houghton L et al (1999) Future of health insurance. N Engl J Med 965:325–329

25. South J, Blass B (2001) The future of modern genomics. Blackwell, London